OpenStudy (anonymous):

I got 0.907 KJ for ΔHcomb but I don't know if its correct. I also don't know how to write the equation. -------- question: A calorimeter is filled with 10.0 mmole of methane gas and an excess of oxygen. When burned, the ignition wire releases 107.2 J of heat. The heat capacity of the calorimeter (including the bomb and water) is 4.319 kJ oC-1. The initial temperature is 24.75 degrees celsius and the final temperature is 26.85 degrees celsius Calculate ΔHcomb for methane and determine the thermochemical reaction?

OpenStudy (abmon98):

10CH4+20O2-->10CO2+20H2O q = mass x specific heat capacity x change in temperature ΔH = q/1000 ÷ moles Number of Moles=Mass(g)/Molar Mass(g/mol)